Younan Corporation manufactures two products: Product E47F and Product D02M. The company is considering implementing an activity-based costing (ABC) system that allocates its manufacturing overhead to four cost pools. The following additional information is available for the company as a whole and for Products E47F and D02M.

Activity Cost Pool Activity Measure Total Cost Total Activity
Machining Machine-hours $ 165,000 11,000 MHs
Machine setups Number of setups $ 150,000 250 setups
Product design Number of products $ 70,000 2 products
Order size Direct labor-hours $ 310,000 10,000 DLHs
Activity Measure Product E47F Product D02M
Machine-hours 5,000 6,000
Number of setups 200 50
Number of products 1 1
Direct labor-hours 4,000 6,000

Using the ABC system, the percentage of the total overhead cost that is assigned to Product E47F is closest to:

rev: 03_17_2018_QC_CS-119201

Multiple Choice

  • 17.84%

  • 33.09%

  • 50.00%

  • 50.94%
Answer #1
6
Total overhead cost = 165000+150000+70000+310000= $695000
Product D02M
Machining 75000 =165000/11000*5000
Machine setups 120000 =150000/250*200
Product design 35000 =70000/2*1
Order size 124000 =310000/10000*4000
Total 354000
Percentage of the total overhead cost that is assigned to Product D02M = 354000/695000= 50.94%